# Leveraging recurrence in neural network wavefunctions for large-scale simulations of Heisenberg antiferromagnets: the square lattice

### v2

@@ Abstract
Machine-learning-based variational Monte Carlo simulations are a promising approach for targeting quantum many-body ground states, especially in two dimensions and in cases where the ground state is known to have a non-trivial sign structure. While many state-of-the-art variational energies have been reached with these methods for finite-size systems, little work has been done to use these results to extract information about the target state in the thermodynamic limit. In this work, we employ recurrent neural networks (RNNs) as a variational ansätze, and leverage their recurrent nature to simulate the ground states of progressively larger systems through iterative retraining. This transfer learning technique allows us to simulate spin-$\frac{1}{2}$ systems on lattices with more than 1,000 spins without beginning optimization from scratch for each system size, thus reducing the demands for computational resources. In this study, we focus on the square-lattice antiferromagnetic Heisenberg model, where it is possible to carefully benchmark our results. We show that we are able to systematically improve the accuracy of the results from our simulations by increasing the training time, and obtain results for finite-sized lattices that are in good agreement with the literature values. Furthermore, we use these results to extract accurate estimates of the ground-state properties in the thermodynamic limit. This work demonstrates that RNN wavefunctions can be used to accurately study quantum many-body physics in the thermodynamic limit.
@@ II.1 The Heisenberg antiferromagnet
where $\langle ij\rangle$ are nearest-neighbor interactions on an $L\times L$ square lattice with $N=L^{2}$ spin-$\frac{1}{2}$ spins. Importantly, the SLAHM can be made stoquastic using the Marshall sign rule [45, 46], which, by the Perron-Frobenius theorem [47, 48] renders the ground-state wavefunction sign-free. As such, there is no sign problem for QMC when the Marshall sign rule is applied to the SLAHM. Here, we also apply the Marshall sign rule so that we can assume a positive variational wavefunction. More details on how the Marshall sign rule is imposed can be found in Appendix A.
@@
In fact, the initial assumption required to apply spin-wave theory to the SLAHM is that the continuous SU(2) symmetry is broken for the ground state in the thermodynamic limit. In his initial calculations, Anderson crucially (and correctly) assumed a “large, constant z-component of spin on one of the two sublattices” [5], thus choosing a direction for the total spin. While this assumption explicitly violates global spin rotation, i.e. the SU(2) symmetry, it can be justified by thinking of the ground state as a superposition, or “wave-packet”, of SU(2)-symmetric states that yields the desired large, constant z-component of the total spin [5, 50, 51]. The energy cost to form such a state decreases as $1/N$, meaning this assumption and subsequent interpretation is valid in the thermodynamic limit where $N\rightarrow\infty$. In other words, there exists a set of states $|0\rangle$, which includes the ground state and some low-lying excited states that collapse onto the ground state in the thermodynamic limit as $1/N$ [5, 52]. This set of states $|0\rangle$ is known as Anderson’s tower of states and describes the spontaneous SU(2) symmetry breaking of the SLAHM. Interestingly, from the symmetry-broken ground state, spin-wave theory correctly predicted the emergence of gapless magnon excitations. These excitations were later identified as the Goldstone modes associated with the continuous symmetry breaking exhibited by the SLAHM [53].
@@ II.2 Variational optimization
As a consequence of the variational principle, the variational energy $E_{\mathcal{W}}$ provides an upper bound to the true ground-state energy. We can thus approximate the true ground state of $\hat{H}$ with our variational ansatz $|\Psi_{\mathcal{W}}\rangle$ by minimizing the variational energy with an appropriately chosen optimization scheme. In this work, we use the Adam optimizer [55] to minimize $E_{\mathcal{W}}$.
@@ II.3 Recurrent neural network wavefunctions
After obtaining all of the conditional probability distributions over each spin according to the predefined autoregressive sequence, one can construct the variational wavefunction as follows: when a ground state has real and positive amplitudes, as is the case for the SLAHM when the Marshall sign rule is applied, the conditional amplitude can be obtained directly from the conditional probability, i.e. $\psi(\sigma_{i}|\sigma_{j<i})=\sqrt{p_{\mathcal{W}}(\sigma_{i}|\sigma_{j<i})}$. The amplitude of the full spin configuration $\vec{\sigma}$ is the product of these local amplitudes $\psi(\vec{\sigma})=\prod_{i}\psi(\sigma_{i}|\sigma_{j<i})$. Appendix B contains a more detailed description of the RNN wavefunctions used in this work.
@@ II.4 Symmetries
Incorporating symmetries in a variational ansatz can dramatically reduce the size of the effective Hilbert space and improve variational accuracies [66, 67]. For the SLAHM, two types of symmetries can be exploited. The first are those that arise due to the Hamiltonian. Efficiently imposing the full SU(2) symmetry of the SLAHM and its eigenstates in NQS is possible by working in the angular momentum basis [68, 63]. However, it is known that the finite-size ground states of the SLAHM are singlet states, which also obey U(1) symmetry [45, 49]. U(1) symmetry can be explicitly enforced in the z-basis by constraining the state to be in the correct magnetization sector [57, 69, 38]. For simplicity, we adopt the latter approach.
@@ II.5 Iterative retraining
It is worth mentioning that even though the number of parameters in an RNN wavefunction with shared weights does not explicitly depend on the number of spins in the system $N$, the hidden dimension $d_{h}$ should display some scaling with $N$ in order to reach sufficiently accurate results. For iterative retraining to be successful, one must choose a sufficiently large $d_{h}$, so that the RNN wavefunction is expressive enough to represent the ground state for the largest system considered. The exact scaling of $d_{h}(N)$ is an open question.
@@ III Results and discussion
We focus on our results for the SLAHM with periodic boundary conditions only, as it is the case which has been studied more thoroughly both theoretically and numerically. We compare against reference values from various QMC studies. The reference ground-state energies for finite-size systems will appear in a forthcoming manuscript [75]. The reference value for the ground-state energy in the thermodynamic limit is from Ref. [9]. All estimates for the squared sublattice magnetization are taken from Ref. [10]. We take these reference values to be “ground truth”. We performed a similar analysis for the SLAHM with open boundary conditions, which both supports and expands on previous RNN results [35, 38]. Those results can be found in Appendix F.
@@ III.1 Energy scaling
The first quantity to assess when using a variational method is the variational energy. When the energy error is much smaller than the energy gap between the ground state and the first excited state, one can safely assume that the largest contribution to the variational wavefunction is the overlap with the true ground-state wavefunction  [54]. Within this regime, the accuracy of observables can be also estimated [76, 54, 77].
@@
Figure 3: (a) The final variational energies obtained from all optimized RNN wavefunctions for each system size plotted according to the known scaling form of the energy, $1/L^{3}$ [78, 52]. All of the final variational energies were estimated with $10\times 10^{3}$ samples. Reference energies for each $L$ [75] and the thermodynamic limit (TL) value of the ground-state energy per spin [9] from QMC is shown for reference. The inset shows a zoomed view of the region close to zero plotted on a log scale for easier viewing. (b) The variances of the above final variational energies as a function of system size $L$.
@@
An interesting trend to note in Fig. 3(a) and (b) is the behavior of the variational energy and variance of the variational energy from a given simulation as a function of system size $L$. As mentioned, we observe that some variational energies in Fig. 3(a) diverge from the reference QMC energies for larger $L$. This result suggests that the variational wavefunction has growing contributions from excited states. However, in all cases, the variances of the variational energies decrease for larger system sizes. We offer one interpretation of this trend for the cases where our variational energies do not diverge from the QMC values. For our simulations performed with the largest scale $s=4.0$, the scaling of the variational energy with system size appears to be well behaved and these variational energies have the lowest variances. It is undeniable that there are contributions to the variational wavefunctions from excited states, as implied by the non-zero variances. It is possible, however, that these contributions come from the low-lying excited states in Anderson’s tower of states $|0\rangle$ which are collapsing onto the ground state as $N\rightarrow\infty$. This offers an explanation for both the improved variational energies and the decreasing variance of the variational energies.
@@
In any case, Fig. 3(a) and (b) show that our results are made systematically better by either increasing the scale $s$ or decreasing the rate $r$ of the training schedule. When one has such a convergent sequence of states, it is possible to extrapolate the variational energy to the zero-variance limit [81, 82, 83, 84, 32, 85]. To obtain estimates of zero-variance energies and appropriate error bars from our variational energies, we first perform a linear fit through the best of the variational energies and their respective variances for a given lattice size $L$. Then, using that linear fit and a wild bootstrap method [86], with $10^{3}$ bootstraps, we obtain more statistically accurate estimates for the zero-variance energies and their error bars. We extrapolated our results from simulations with the three largest scales $s=1.0,2.0,4.0$, which produced the most accurate variational energies, as seen in the inset of Fig. 3(a), as well as the smallest variances on the variational energies, as seen in Fig. 3(b). The resulting zero-variance energies, shown in Fig. 4, are in close agreement with the reference energies from QMC.
@@ III.2 Sublattice magnetization scaling
where $\vec{r}=|\vec{r}_{i}-\vec{r}_{j}|$. The $(z)$ indicates which real space correlations defined in Eq. 8 or Eq. 9 were used. Of particular interest is the staggered structure factor, which is defined as $S^{(z)}\left(L,\vec{q}=(\pi,\pi)\right)$.
@@
Figure 5: Estimates of the squared sublattice magnetization scaled as a function of $1/L$. Estimates of $M^{2}$ and $M_{C}^{2}$ are shown to demonstrate that they should extrapolate to the same value in the thermodynamic limit. The inset shows a zoomed view of the y-intercept. Closed markers represent the values of $M^{2}$ and $M_{C}^{2}$ estimated using the correlations defined by Eq. 8, while open markers represent the values of $M^{2}$ and $M^{2}_{C}$ that are estimated using the correlations defined by Eq. 9. Finite-size estimates and the thermodynamic limit (TL) value of the squared sublattice magnetization from QMC [10] are shown for reference.
Fig. 5 shows the estimates of Eq. 11 and Eq. 12 obtained from our most accurate simulations (scale $s=4.0$). Each real-space correlation was estimated with $10\times 10^{3}$ samples. Our finite-size estimates of $M^{2}$ and $M^{2}_{C}$ appear to be in close agreement to the available QMC reference values up to very large system size $L$; however, our estimates of $M^{2}_{z}$ and $M_{C^{z}}^{2}$ do not exhibit the expected behavior. The growing disagreement between $M^{2}$ and $M^{2}_{z}$ (or equivalently between $M^{2}_{C}$ and $M^{2}_{C^{z}}$) indicates that the SU(2) symmetry of our RNN wavefunction is broken during the iterative retraining. Recalling that the existence of Anderson’s tower of states $|0\rangle$ offers a mechanism to explain the spontaneous SU(2) symmetry breaking in the SLAHM [5, 51], we conclude that the observed symmetry breaking in our RNN wavefunctions supports the hypothesis put forward in Section III.1 that the states learned during our best simulations ($s=4.0$) contain only contributions from the low-lying excited states in $|0\rangle$.
@@
The symmetry breaking behavior is striking for both rates shown in Fig. 5, but we notice that the divergence between e.g. $M^{2}$ and $M^{2}_{z}$ for $r=0.25$ is slower than the divergence for $r=0.475$. In general, we notice that for larger scales or slower rates (smaller r), the symmetry breaking occurs later and less dramatically, suggesting that longer training time might lead our RNN wavefunctions to restore the broken SU(2) symmetry. Interestingly, we only observe this symmetry breaking in our simulations for systems with periodic boundary conditions (see Fig. 13 in Appendix F). It is possible that this behavior is the result of the expressiveness (or lack thereof) of the recurrent cell employed for the periodic RNN wavefunctions (see Appendix B). To this point, some studies have shown that broken symmetries in an NQS can be restored by increasing the expressiveness of the ansatz [34].
@@
Nevertheless, we are able to extract accurate estimates of the sublattice magnetization in the thermodynamic limit from our simulations. Indeed, this is because the lowest-lying states in Anderson’s tower of states exhibit the same squared value of the sublattice magnetization, i.e. the same values for $M^{2}_{(z)}$ and $M^{2}_{C^{(z)}}$ [50, 87, 51]. To obtain our extrapolated values of the sublattice magnetization, we perform multiple polynomial fits including up to the third order in $1/L$. Sandvik et. al [9] performed a constrained third-order polynomial regression and argued that the additional sub-leading terms were essential for obtaining an accurate estimate of the extrapolated value. In  Fig. 5, we show only the linear and second-order polynomial fits because we find that fitting a third-order polynomial to our data leads to spurious behavior outside of the data range.
@@
Fitting a second-order polynomial to $M^{2}$ yields extrapolated values for $M_{\infty}$ that are within the error bars of $M^{C}_{\infty}$, the values obtained from performing either a first- or second-order fit on $M^{2}_{c}$. The agreement between these extrapolated values can be seen in the inset of Fig. 5. The linear fit on $M^{2}_{c}$ provides tighter error bars, so we report $M^{C}_{\infty}$ obtained from that fit. These estimates for $M^{2}$ and $M^{2}_{C}$ are not only in good agreement with each other, but also with the reference value obtained from QMC simulations $M_{\infty}^{\text{QMC}}=0.30743(1)$ [10]. We obtain extrapolated values for the sublattice magnetization of $M_{\infty}=0.309(2)$ and $M_{\infty}^{C}=0.308(2)$ using a second-order polynomial fit on our estimates of $M^{2}$ and a linear fit on our estimates of $M^{2}_{C}$ and then improving our statistical estimates with a wild bootstrap [86] method and $10\times 10^{3}$ bootstraps.

### v3

# Leveraging recurrence in neural network wavefunctions for large-scale simulations of Heisenberg antiferromagnets on the square lattice

@@ Abstract
Machine-learning-based variational Monte Carlo simulations are a promising approach for targeting quantum many-body ground states, especially in two dimensions and in cases where the ground state is known to have a non-trivial sign structure. While many state-of-the-art variational energies have been reached with these methods for finite-size systems, little work has been done to use these results to extract information about the target state in the thermodynamic limit. In this work, we employ recurrent neural networks (RNNs) as a variational ansätze, and leverage their recurrent nature to simulate the ground states of progressively larger systems through iterative retraining. This transfer learning technique allows us to simulate spin-$\frac{1}{2}$ systems on lattices with more than 1,000 spins without beginning optimization from scratch for each system size, thus reducing the demands for computational resources. In this study, we focus on the square-lattice antiferromagnetic Heisenberg model, where it is possible to carefully benchmark our results. We show that we are able to systematically improve the accuracy of the results from our simulations by increasing the training time, and obtain results for finite-sized lattices that are in good agreement with the literature values. Furthermore, we use these results to extract accurate estimates of the ground-state properties in the thermodynamic limit. This work demonstrates that RNN wavefunctions are able to extract accurate information about quantum many-body systems in the thermodynamic limit.
@@ II.1 The Heisenberg antiferromagnet
where $\langle ij\rangle$ are nearest-neighbor interactions on an $L\times L$ square lattice with $N=L^{2}$ $\frac{1}{2} spin$s. Importantly, the SLAHM can be made stoquastic using the Marshall-Peierls sign rule [45, 46], which, by the Perron-Frobenius theorem [47, 48] renders the ground-state wavefunction sign-free. As such, there is no sign problem for QMC when the Marshall-Peierls sign rule is applied to the SLAHM. Here, we also apply the Marshall-Peierls sign rule so that we can assume a positive variational wavefunction. More details on how the Marshall-Peierls sign rule is imposed can be found in Appendix A.
@@
In fact, the initial assumption required to apply spin-wave theory to the SLAHM is that the continuous SU(2) symmetry is broken for the ground state in the thermodynamic limit. In his initial calculations, Anderson crucially (and correctly) assumed a “large, constant $z$ component of spin on one of the two sublattices” [5], thus choosing a direction for the total spin. While this assumption explicitly violates global spin rotation, i.e. the SU(2) symmetry, it can be justified by thinking of the ground state as a superposition, or “wave-packet”, of SU(2)-symmetric states that yields the desired large, constant $z$ component of the total spin [5, 50, 51]. The energy cost to form such a state decreases as $1/N$, meaning this assumption and subsequent interpretation is valid in the thermodynamic limit where $N\rightarrow\infty$. In other words, there exists a set of states $|0\rangle$, which includes the ground state and some low-lying excited states that collapse onto the ground state in the thermodynamic limit as $1/N$ [5, 52]. This set of states $|0\rangle$ is known as Anderson’s tower of states and describes the spontaneous SU(2) symmetry breaking of the SLAHM. Interestingly, from the symmetry-broken ground state, spin-wave theory correctly predicted the emergence of gapless magnon excitations. These excitations were later identified as the Goldstone modes associated with the continuous symmetry breaking exhibited by the SLAHM [53].
@@ II.2 Variational optimization
As a consequence of the variational principle, the variational energy $E_{\mathcal{W}}$ provides an upper bound to the true ground-state energy. We can thus approximate the true ground state of $\hat{H}$ with our variational ansatz $|\Psi_{\mathcal{W}}\rangle$ by minimizing the variational energy with an appropriately chosen optimization scheme [54, 55]. In this work, we use the Adam optimizer [56] to minimize $E_{\mathcal{W}}$.
@@ II.3 Recurrent neural network wavefunctions
After obtaining all of the conditional probability distributions over each spin according to the predefined autoregressive sequence, one can construct the variational wavefunction as follows: when a ground state has real and positive amplitudes, as is the case for the SLAHM when the Marshall-Peierls sign rule is applied, the conditional amplitude can be obtained directly from the conditional probability, i.e. $\psi(\sigma_{i}|\sigma_{j<i})=\sqrt{p_{\mathcal{W}}(\sigma_{i}|\sigma_{j<i})}$. The amplitude of the full spin configuration $\vec{\sigma}$ is the product of these local amplitudes $\psi(\vec{\sigma})=\prod_{i}\psi(\sigma_{i}|\sigma_{j<i})$. Appendix B contains a more detailed description of the RNN wavefunctions used in this work.
@@ II.4 Symmetries
Incorporating symmetries in a variational ansatz can dramatically reduce the size of the effective Hilbert space and improve variational accuracies [66, 67]. For the SLAHM, two types of symmetries can be exploited. The first are those that arise due to the Hamiltonian. Efficiently imposing the full SU(2) symmetry of the SLAHM and its eigenstates in NQS is possible by working in the angular momentum basis [68, 63]. However, it is known that the finite-size ground states of the SLAHM are singlet states, which also obey U(1) symmetry [45, 49]. U(1) symmetry can be explicitly enforced in the $z$ basis by constraining the state to be in the correct magnetization sector [55, 69, 38]. For simplicity, we adopt the latter approach.
@@ II.5 Iterative retraining
It is worth mentioning that even though the number of parameters in an RNN wavefunction with shared weights does not explicitly depend on the number of spins in the system $N$, the hidden dimension $d_{h}$ should display some scaling with $N$ in order to reach sufficiently accurate results. For iterative retraining to be successful, one must choose a sufficiently large $d_{h}$, so that the RNN wavefunction is expressive enough to represent the ground state for the largest system considered. The exact scaling of $d_{h}(N)$ is an open question, though it has been studied for other architectures [75].
@@ III Results and discussion
We focus on our results for the SLAHM with periodic boundary conditions only, as it is the case which has been studied more thoroughly both theoretically and numerically. We compare against reference values from various QMC studies. The reference ground-state energies for finite-size systems will appear in a forthcoming manuscript [76]. The reference value for the ground-state energy in the thermodynamic limit is from Ref. [9]. All estimates for the squared sublattice magnetization are taken from Ref. [10]. We take these reference values to be “ground truth”. We performed a similar analysis for the SLAHM with open boundary conditions, which both supports and expands on previous RNN results [35, 38]. Those results can be found in Appendix G.
@@ III.1 Energy scaling
The first quantity to assess when using a variational method is the variational energy. When the energy error is much smaller than the energy gap between the ground state and the first excited state, one can safely assume that the largest contribution to the variational wavefunction is the overlap with the true ground-state wavefunction  [54]. Within this regime, the accuracy of observables can be also estimated [54, 77].
@@
Figure 3: (a) The final variational energies obtained from all optimized RNN wavefunctions for each system size plotted according to the known scaling form of the energy, $1/L^{3}$ [78, 52]. All of the final variational energies were estimated with $10\times 10^{3}$ samples. Reference energies for each $L$ [76] and the thermodynamic limit (TL) value of the ground-state energy per spin [9] from QMC is shown for reference. The inset shows a zoomed view of the region close to zero plotted on a log scale for easier viewing. (b) The variances of the above final variational energies as a function of the number of spins in the system $N$.
@@
Fig. 3(a) and (b) show that our results are made systematically better by either increasing the scale $s$ or decreasing the rate $r$ of the training schedule. When one has such a convergent sequence of states, it is possible to extrapolate the variational energy to the zero-variance limit [81, 82, 83, 84, 32, 85]. To obtain estimates of zero-variance energies and appropriate error bars from our variational energies, we first perform a linear fit through the best of the variational energies and their respective variances for a given lattice size $L$  [^1] We extrapolated our results from simulations with the three largest scales $s=1.0,2.0,4.0$, which produced the most accurate variational energies, as seen in the inset of Fig. 3(a), as well as the smallest variances on the variational energies, as seen in Fig. 3(b).. See Appendix F for an example of our zero-variance extrapolation for $L=32$. Using that linear fit and a wild bootstrap method [87], with $10^{3}$ bootstraps, we obtain more statistically accurate estimates for the zero-variance energies and their error bars. The resulting zero-variance energies, shown in Fig. 4, are in close agreement with the reference energies from QMC.
@@ III.2 Sublattice magnetization scaling
where $\vec{r}=|\vec{r}_{i}-\vec{r}_{j}|$. The superscript $(z)$ indicates which real space correlations defined in Eq. 8 or Eq. 9 were used. Of particular interest is the staggered structure factor, which is defined as $S^{(z)}\left(L,\vec{q}=(\pi,\pi)\right)$.
@@
Figure 5: Estimates of the squared sublattice magnetization scaled as a function of $1/L$. Estimates of $M^{2}$ and $M_{C}^{2}$ are shown to demonstrate that they should extrapolate to the same value in the thermodynamic limit. The inset shows a zoomed view of the $y$ intercept. Closed markers represent the values of $M^{2}$ and $M_{C}^{2}$ estimated using the correlations defined by Eq. 8, while open markers represent the values of $M^{2}$ and $M^{2}_{C}$ that are estimated using the correlations defined by Eq. 9. Finite-size estimates and the thermodynamic limit (TL) value of the squared sublattice magnetization from QMC [10] are shown for reference.
Fig. 5 shows the estimates of Eq. 11 and Eq. 12 obtained from our most accurate simulations (scale $s=4.0$). Each real-space correlation was estimated with $10\times 10^{3}$ samples. We only calculate $M^{2}$ for $L\leq 20$ because the calculation scales as $\mathcal{O}(N^{3})$: Eq. 10 involves a sum over all two-point correlations and performing inference with our RNN wavefunction scales as $\mathcal{O}(N)$.
@@
Our finite-size estimates of $M^{2}$ and $M^{2}_{C}$ appear to be in close agreement to the available QMC reference values up to very large system size $L$; however, our estimates of $M^{2}_{z}$ and $M_{C^{z}}^{2}$ do not exhibit the expected behavior. The growing disagreement between $M^{2}$ and $M^{2}_{z}$ (or equivalently between $M^{2}_{C}$ and $M^{2}_{C^{z}}$) indicates that the SU(2) symmetry of our RNN wavefunction is broken during the iterative retraining. The symmetry breaking behavior is striking for both rates shown in Fig. 5, but we notice that the divergence between e.g. $M^{2}$ and $M^{2}_{z}$ for $r=0.25$ is slower than the divergence for $r=0.475$. In general, we notice that for larger scales or slower rates (smaller r), the symmetry breaking occurs later and less dramatically. It is possible that even longer training time might allow our RNN wavefunctions to restore the broken SU(2) symmetry. Interestingly, we only observe this symmetry breaking in our simulations for systems with periodic boundary conditions (see Fig. 14 in Appendix G). It is possible that this behavior is the result of the expressiveness (or lack thereof) of the recurrent cell employed for the periodic RNN wavefunctions (see Appendix B). To this point, some studies have shown that broken symmetries in an NQS can be restored by increasing the expressiveness of the ansatz [34].
@@
Despite the observed symmetry breaking, we are still able to extract accurate estimates of the sublattice magnetization in the thermodynamic limit from the estimates of $M^{2}$ and $M^{2}_{C}$ from our simulations. To obtain our extrapolated values of the sublattice magnetization, we perform multiple polynomial fits including up to the third order in $1/L$. Sandvik et. al [9] performed a constrained third-order polynomial regression and argued that the additional sub-leading terms were essential for obtaining an accurate estimate of the extrapolated value. In  Fig. 5, we show only the linear and second-order polynomial fits because we find that fitting a third-order polynomial to our data leads to spurious behavior outside of the data range.
@@
Fitting a second-order polynomial to $M^{2}$ yields extrapolated values for $M_{\infty}$ that are within the error bars of $M^{C}_{\infty}$, the values obtained from performing either a first- or second-order fit on $M^{2}_{c}$. The agreement between these extrapolated values can be seen in the inset of Fig. 5. The linear fit on $M^{2}_{c}$ provides tighter error bars, so we report $M^{C}_{\infty}$ obtained from that fit. Our extrapolated estimates for $M^{2}$ and $M^{2}_{C}$ are not only in good agreement with each other, but also with the reference value obtained from QMC simulations $M_{\infty}^{\text{QMC}}=0.30743(1)$ [10]. We obtain extrapolated values for the sublattice magnetization of $M_{\infty}=0.309(2)$ and $M_{\infty}^{C}=0.308(2)$ using a second-order polynomial fit on our estimates of $M^{2}$ and a linear fit on our estimates of $M^{2}_{C}$ and then improving our statistical estimates with a wild bootstrap [87] method and $10\times 10^{3}$ bootstraps.
Our results for the sublattice magnetization strongly suggest that the states learned by our RNN wavefunctions are a superposition of the states in Anderson’s tower of states $|0\rangle$. Anderson’s tower of states was an idea originally proposed to offer a mechanism that explains the spontaneous SU(2) symmetry breaking in the SLAHM [5, 51]. Importantly, all of the lowest-lying states in $|0\rangle$ exhibit the same squared value of the sublattice magnetization, i.e. the same values for $M^{2}$, $M^{2}_{z}$, $M^{2}_{C}$, and $M^{2}_{C^{z}}$ [50, 88, 51]. The observed symmetry breaking in our RNN wavefunctions means that our wavefunctions contain contributions from states other than the true finite-size ground state, which is an SU(2)-symmetric singlet state. This is consistent with the fact that the variances of our variational energies are non-zero. However, our ability to extract accurate estimates of $M_{\infty}$ and $M_{\infty}^{C}$ suggests that the extra contributions are dominated by the states in $|0\rangle$.